# A probabilistic, data-driven closure model for RANS simulations with aleatoric, model uncertainty

### v1

@@ 4 Conclusions
In order to enable inference and learning tasks, we developed a differentiable RANS solver capable of providing parametric sensitivities. Such a differentiable solver was non-trivial owing to the complexity of the physical simulator and its stability issues. The lack of such numerical tools has proven to be a significant barrier for intrusive, physics-based, data-driven models in turbulence Cranmer et al. (2020). This differentiable solver was utilized in the context of a Stochastic Variational Inference (SVI) scheme that employs Monte Carlo estimates of the ELBO derivatives in conjunction with the reparametrization trick and stochastic gradient ascent. We demonstrated how probabilistic predictive estimates can be computed for all output quantities of the trained RANS model and illustrated their accuracy on a separated flow in the backward facing step benchmark problem. In most cases very good agreement with the reference values was achieved and in all cases these were enveloped by the credible intervals computed.
@@
  • The parametric closure model could benefit from non-local dependencies which could be enabled by convolutional or vector-cloud neural networks (VCNN) Han et al. (2022) with appropriate embedding of invariance properties.
  • The dimensionality reduction of the stochastic discrepancy terms (Equation (23)) was based on a pre-selected and uniform division of the problem domain into subdomains. The accuracy of the model would certainly benefit from a learnable and adaptive such scheme that would be able to focus on the areas where model deficiencies are most pronounced and stochastic corrections are most needed.
@@ Appendix A Differentiable RANS solver
In the present study, the RANS equations (Equation (7)) are numerically solved using the finite element discretization, implemented in the open source package FEniCS Alnæs et al. (2015). The discrete equations are obtained by representing the solution and test functions in appropriate finite dimensional function spaces. In particular, we employed the standard Taylor-Hood pair of basis functions with polynomial degree one for the pressure interpolants and two for the velocities. This choice is made to avoid stability issues potentially arising from the interaction between the momentum and continuity equations.
The turbulence scaling terms, $k$ and $\epsilon$ in Equation (21), are obtained by solving the respective standard transport equations Pope (2000); Alfonsi (2009). Symmetry is enforced in the RS tensor, i.e. $\tau_{xy}$ and $\tau_{yx}$ are identical without any redundancy in the representation. The discretized system is solved with damped Newton’s method. For robustness and global convergence, pseudo-time stepping is used with the backward Euler discretization Deuflhard (2005). As the Reynolds number is increased, the convection term dominates, leading to stability Donea and Huerta (2003).This elicits a need to add stabilization terms to the weak form, such as the least-square stabilization, according to which the weighted square of the strong form is added to the weak form residual. However, these extra terms have to be chosen carefully in order not to compromise the correctness of the approximate solution. Classically, researchers added artificial diffusion terms or a numerical diffusion by using upwind scheme for the convection term instead of central diffusion. The extra infused term corrupted the solution quality. To avoid this, in practice, it is common to use schemes like Streamline-Upwind Petrov-Gelarkin method (SUPG) and Galerkin Least Squares (GLS). In the present study, we have utilized a self-adjoint numerical stabilisation scheme which is an extension of Gelarkin Least Squares (GLS) Stabilisation called Galerkin gradient least square method Franca and Do Carmo (1989). This amounts to adding a stabilization term to the residual weak form. For additional details, interested readers are referred to Franca and Do Carmo (1989); Donea and Huerta (2003).
@@ Appendix B Adjoint Formulation and Estimation of the Gradient of the ELBO
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})=\mathbb{E}_{q(\bm{\phi})}\left[\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta})-\log q_{\xi}(\bm{\eta})\right)\right]$ (38)
@@
$\begin{array}[]{ll}\log p(\mathcal{D},\bm{\eta})&=\log\left(p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})\right)\\ &=\left(\sum_{m=1}^{M}\log p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{E}_{\tau}^{(m)})+\log p(\bm{E}_{\tau}^{(m)}|\bm{\Lambda})\right)~{}+\log p(\bm{\theta})+\log~{}p(\bm{\Lambda})\end{array}$ (39)
@@
In particular, and if we drop the superscript $m$ for each term in the log-likelihood in order to simplify the notation, we formulate a Lagrangian with the help of a vector $\lambda$ of Lagrangian multipliers, i.e.:
$\mathcal{L}=\ell+\bm{\lambda}^{T}(\mathcal{G}(\bm{z})-\bm{B}\bm{\tau})$ (41)
@@
$=\left(\frac{\partial\ell}{\partial\bm{z}}+\bm{\lambda}^{T}\frac{\partial\mathcal{G}}{\partial\bm{z}}\right)\frac{d\bm{z}}{d\bm{\bm{\tau}}}-\bm{\lambda}^{T}\bm{B}$ (42)
@@
$\frac{\partial\ell}{\partial\bm{z}}+\bm{\lambda}^{T}\frac{\partial\mathcal{G}}{\partial\bm{z}}=0\quad\textrm{or,}\quad\left(\frac{\partial\mathcal{G}}{\partial\bm{z}}\right)^{T}\bm{\lambda}=-\left(\frac{\partial\ell}{\partial\bm{z}}\right)^{T}$ (43)
@@
$\frac{d\mathcal{L}}{d\bm{\tau}}=\frac{d\ell}{d\bm{\tau}}=-\bm{\lambda}^{T}\bm{B}$ (44)
@@
$\frac{d\ell}{d\bm{\theta}}=\underbrace{\frac{\partial\ell}{\partial\bm{\tau}}}_{\begin{subarray}{c}\text{Adjoint}\\ \text{model}\end{subarray}}\underbrace{\frac{\partial\bm{\tau}}{\partial\bm{\theta}}}_{\begin{subarray}{c}\text{NN}\\ \text{auto-diff}\end{subarray}}$ (49)
@@
$\mathcal{F}(\bm{\xi})\approx\frac{1}{K}\left(\sum_{k=1}^{K}\left(\sum_{m=1}^{M}\ell^{(m)}(\bm{\theta},\bm{E}_{\tau}^{(m,k)})+\log p(\bm{E}_{\tau}^{(m,k)}|\bm{\Lambda})\right)~{}+\log p(\bm{\theta})+\log~{}p(\bm{\Lambda})-\log q_{\xi}(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(m,k)})\right)$ (51)
@@
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})\approx\frac{1}{K}\sum_{k=1}^{K}\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi}^{(k)})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta}^{(k)})-\log q_{\xi}(\bm{\eta}^{(k)})\right)\$ (53)

### v2

@@ 4 Conclusions
In order to enable inference and learning tasks, we developed a differentiable RANS solver capable of providing parametric sensitivities. Such a differentiable solver was non-trivial owing to the complexity of the physical simulator and its stability issues. The lack of such numerical tools has proven to be a significant barrier for intrusive, physics-based, data-driven models in turbulence [26]. This differentiable solver was utilized in the context of a Stochastic Variational Inference (SVI) scheme that employs Monte Carlo estimates of the ELBO derivatives in conjunction with the reparametrization trick and stochastic gradient ascent. We demonstrated how probabilistic predictive estimates can be computed for all output quantities of the trained RANS model and illustrated their accuracy on a separated flow in the backward facing step benchmark problem. In most cases, very good agreement with the reference values was achieved and in all cases these were enveloped by the credible intervals computed.
@@
  • The parametric closure model could benefit from non-local dependencies which could be enabled by convolutional or vector-cloud neural networks (VCNN) [93] with appropriate embedding of invariance properties.
  • The dimensionality reduction of the stochastic discrepancy terms (Equation (19)) was based on a pre-selected and uniform division of the problem domain into subdomains. The accuracy of the model would certainly benefit from a learnable and adaptive scheme that would be able to focus on the areas where model deficiencies are most pronounced and stochastic corrections are most needed.
Acknowledgements
We extend our sincere thanks to Jigar Parekh for his valuable guidance in the generation of the LES data and to Maximilian Rixner for his assistance in implementing the forward solver. We are also thankful to Nicholas Geneva and Niklas Fehn for their valuable discussions that enriched this work. Our sincere gratitude is extended to the anonymous reviewers, whose insightful feedback greatly enhanced the quality of our manuscript.
@@ Appendix A Differentiable RANS solver
In the present study, the RANS equations (Equation (7)) are numerically solved using the finite element discretization, implemented in the open source package FEniCS [87]. The discrete equations are obtained by representing the solution and test functions in appropriate finite dimensional function spaces. In particular, we employed the standard Taylor-Hood pair of basis functions with polynomial degree one for the pressure interpolants and two for the velocities. This choice is made to avoid stability issues potentially arising from the interaction between the momentum and continuity equations.
The turbulence scaling terms, $k$ and $\epsilon$ in Equation (17), are obtained by solving the respective standard transport equations [2, 52]. Symmetry is enforced in the RS tensor, i.e. $\tau_{xy}$ and $\tau_{yx}$ are identical without any redundancy in the representation. The discretized system is solved with damped Newton’s method. For robustness and global convergence, pseudo-time stepping is used with the backward Euler discretization [94]. As the Reynolds number is increased, the convection term dominates, leading to stability [95].This elicits a need to add stabilization terms to the weak form, such as the least-square stabilization, according to which the weighted square of the strong form is added to the weak form residual. However, these extra terms have to be chosen carefully in order not to compromise the correctness of the approximate solution. Classically, researchers added artificial diffusion terms or a numerical diffusion by using upwind scheme for the convection term instead of central diffusion. The extra infused term corrupted the solution quality. To avoid this, in practice, it is common to use schemes like Streamline-Upwind Petrov-Galerkin method (SUPG) and Galerkin Least Squares (GLS). In the present study, we have utilized a self-adjoint numerical stabilisation scheme which is an extension of Galerkin Least Squares (GLS) Stabilisation called Galerkin gradient least square method [96]. This amounts to adding a stabilization term to the residual weak form. For additional details, interested readers are referred to [96, 95].
@@ Appendix B Adjoint Formulation and Estimation of the Gradient of the ELBO
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})=\mathbb{E}_{q(\bm{\phi})}\left[\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta})-\log q_{\xi}(\bm{\eta})\right)\right],$ (34)
@@
$\begin{array}[]{ll}\log p(\mathcal{D},\bm{\eta})&=\log\left(p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})\right)\\ &=\left(\sum_{m=1}^{M}\log p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{E}_{\tau}^{(m)})+\log p(\bm{E}_{\tau}^{(m)}|\bm{\Lambda})\right)\\ &~{}+\log p(\bm{\theta})+\log~{}p(\bm{\Lambda}).\end{array}$ (35)
@@
In particular, and if we drop the superscript $m$ for each term in the log-likelihood in order to simplify the notation, we formulate a Lagrangian with the help of a vector $\bm{\lambda}$ of Lagrangian multipliers, i.e.:
$\mathcal{L}=\ell+\bm{\lambda}^{T}(\mathcal{G}(\bm{z})-\bm{B}\bm{\tau}),$ (37)
@@
$=\left(\frac{\partial\ell}{\partial\bm{z}}+\bm{\lambda}^{T}\frac{\partial\mathcal{G}}{\partial\bm{z}}\right)\frac{d\bm{z}}{d\bm{\bm{\tau}}}-\bm{\lambda}^{T}\bm{B}.$ (38)
@@
$\frac{\partial\ell}{\partial\bm{z}}+\bm{\lambda}^{T}\frac{\partial\mathcal{G}}{\partial\bm{z}}=0\quad\textrm{or,}\quad\left(\frac{\partial\mathcal{G}}{\partial\bm{z}}\right)^{T}\bm{\lambda}=-\left(\frac{\partial\ell}{\partial\bm{z}}\right)^{T}.$ (39)
@@
$\frac{d\mathcal{L}}{d\bm{\tau}}=\frac{d\ell}{d\bm{\tau}}=-\bm{\lambda}^{T}\bm{B}.$ (40)
@@
$\frac{d\ell}{d\bm{\theta}}=\underbrace{\frac{\partial\ell}{\partial\bm{\tau}}}_{\begin{subarray}{c}\text{Adjoint}\\ \text{model}\end{subarray}}\underbrace{\frac{\partial\bm{\tau}}{\partial\bm{\theta}}}_{\begin{subarray}{c}\text{NN}\\ \text{auto-diff}\end{subarray}},$ (41)
@@
$\mathcal{F}(\bm{\xi})\approx\frac{1}{K}\left(\sum_{k=1}^{K}\left(\sum_{m=1}^{M}\ell^{(m)}(\bm{\theta},\bm{E}_{\tau}^{(m,k)})+\log p(\bm{E}_{\tau}^{(m,k)}|\bm{\Lambda})\right)~{}+\log p(\bm{\theta})+\log~{}p(\bm{\Lambda})-\log q_{\bm{\xi}}(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(m,k)})\right),$ (43)
@@
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})\approx\frac{1}{K}\sum_{k=1}^{K}\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi}^{(k)})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta}^{(k)})-\log q_{\xi}(\bm{\eta}^{(k)})\right),\$ (45)